OpenStudy (anonymous):

if 2x+y=20. what is the value of 6x+3y?

OpenStudy (anonymous):

2x+y=20 -y -y 2x=20 x=10 6(10)+3y-3y 60=-3y -20=y I tried

OpenStudy (dls):

2x+y=20 6x+3y=? Simply multiply first equation by 3.

OpenStudy (mathstudent55):

Notice that 6x + 3y = 3(2x + y) You are given that 3x + y = 20, so just substitute 20 for 3x + y.

OpenStudy (cggurumanjunath):

2x+y=20 6x+3y=? Simply multiply first equation by 3.

OpenStudy (cggurumanjunath):